OpenStudy (anonymous):

Find the Mistake: provided below is a solution to the equation 4m − 8 = 2m. Choose the statement that is true about the given solution. Line # 4m − 8 = 2m 1 −2m − 8 = 0 2 −2m = 8 3 m = −4

OpenStudy (anonymous):

4m-8=2m subtracting 2m from both sides 4m-2m-8=0 simplifying 2m-8=0 so option 1 and 2 are not correct 2m-8=0 adding 8 to both sides 2m=8 dividing both sides by 2 m=4